The cumulative number of car accidents from 2000 to 2010 can be modeled by the quadratic expression C = -43x² + 1376x + 12040 where x=0 corresponds to 2000, x=1 corresponds to 2001 and so on until x=10 corresponds to 2010. Find the number of cumulative car accidents in 2005.
I don't understand this problem at all, could someone give me a hand, please?

Answer by Edwin McCravy(20054)

The cumulative number of car accidents from 2000 to 2010 can be modeled by the quadratic expression C = -43x² + 1376x + 12040 where x=0 corresponds to 2000, x=1 corresponds to 2001 and so on until x=10 corresponds to 2010. Find the number of cumulative car accidents in 2005.

If you substitute x = 0 in C = -43x² + 1376x + 12040 you get the
number of car accidents in the year 2000:

C = -43x² + 1376x + 12040
C = -43(0)² + 1376(0) + 12040
C = 0 + 0 + 12040
C = 12040 accidents in 2000

If you substitute x = 1 in C = -43x² + 1376x + 12040 you get the
number of car accidents in the year 2001:

C = -43x² + 1376x + 12040
C = -43(1)² + 1376(1) + 12040
C = -43 + 1376 + 12040
C = 13373 accidents in 2001

If you substitute x = 2 in C = -43x² + 1376x + 12040 you get the
number of car accidents in the year 2002:

C = -43x² + 1376x + 12040
C = -43(2)² + 1376(2) + 12040
C = -43*4 + 2752 + 12040
C = -52 + 2752 + 12040
C = 14620 accidents in 2002

If you substitute x = 3 in C = -43x² + 1376x + 12040 you get the
number of car accidents in the year 2003:

C = -43x² + 1376x + 12040
C = -43(3)² + 1376(3) + 12040
C = -43*9 + 4128 + 12040
C = -387 + 4128 + 12040
C = 15781 accidents in 2003

If you substitute x = 4 in C = -43x² + 1376x + 12040 you get the
number of car accidents in the year 2004:

C = -43x² + 1376x + 12040
C = -43(4)² + 1376(4) + 12040
C = -43*16 + 5504 + 12040
C = -688 + 5504 + 12040
C = 16856 accidents in 2004

I'll stop there and let you figure out the number of accidents
in the year 2005.

Edwin